OpenStudy (matimaticas):

In the expansion of (3a + 4b)^8, which of the following are possible variable terms? Explain your reasoning. a^2b^3; a^5b^3; ab^8; b^8; a^4b^4; a^8; ab^7; a^6b^5

ganeshie8 (ganeshie8):

hint : the degree of each term has to be `8`

OpenStudy (nirmalnema):

\[(a + b)8 = a^8 + 8a^7b + 28a^6b^2 + 56a^5b^3 + 70a^4b^4 + 56a^3b^5 + 28a^2b^6 + 8ab^7 + b^8.\]

ganeshie8 (ganeshie8):

for example, `a^2b^3` has degree = `2+3` = 5, which is NOT 8. So this is impossible

OpenStudy (matimaticas):

so each term that doesn't have a degree of 8 is a variable term?

ganeshie8 (ganeshie8):

each term that doesn't have a degree of 8 can NOT be a term in the expansion.

OpenStudy (matimaticas):

ok, im just a little confused about what the "variable terms" would mean....